OpenStudy (emetuo89):

a frog is at the bottom of a 18 foot well. each day it craws up 3 feet, but each night it slips back 1 foot. after how many days will the frog reach the top of the well?

OpenStudy (emetuo89):

I'm thinking the answer is 9. (18/3-1) = 18/2 = 9 ?

OpenStudy (anonymous):

Looks fine to me. Since it crawls up 3 ft but falls back 1 ft every night, it means that it crawls up 2 ft each day, essentially.